# Inverse Garment and Pattern Modeling with a Differentiable Simulator

### 11 Mar 2024 v1

@@ 1 Introduction
We evaluate our approach on a variety of garment types and demonstrate that our method produces patterns of promising quality. We compared it against three state-of-the-art comparable to ours, and report that our method can replicate more detailed deformations and produce more accurate pattern estimations. In summary, we make the following contributions:
@@ 2.1 Garment shape recovery from 2D/3D data
Our work is similar in spirit to that of Yang et al [59], who use search-based optimization to recover both 2D sewing patterns (parameterized by numerical, such as sleeve length, waist width) and the 3D garment such that it minimizes the 2D silhouette difference between the projected 3D garment and the input image. The draped 3D garment shape is obtained by using a physics-based simulator [39]. Unlike theirs, in our work the estimation of pattern parameters are coupled with the physically-based simulation process, allowing for direct and precise mapping of 3D garment error to both the pattern geometry and material parameters.
@@ 2.2 Sewing pattern estimation from 3D data
3D-to-2D surface flattening. Several works have considered the garment mesh as developable surface and obtained 2D pattern panels by cutting the garment surface into 3D patches and flattening each of them onto a plane. The cutting lines are found either by projecting the predefined seam lines from the body mesh onto the garment mesh [3], along curvature directions expressed in the cross fields [53, 42] on the surface, or through variational surface cutting[48] to minimize the distortion induced by cutting and flattening [57]. While intuitive and versatile, such surface flattening based on a geometric strategy is prone to generating patterns that deviate significantly from traditional panel semantics, or lack of left-to-right symmetry, making them unsuitable for garment production. Moreover, purely geometric methods [3, 42] do not account for the fabric’s elasticity in physical body-cloth interaction during the draping process, often resulting in a sewing pattern that cannot accurately replicate the originally designed garment.
@@ 2.3 Differentiable cloth simulator
Our method mainly solves an optimization of parameters coupled with the garment simulation process. We follow the thread of differentiable physics simulation [32, 22, 31, 28] that can optimize parameters to fit the measurement data. To our knowledge, this is the first work that applies the differentiable simulator to recover sewing patterns of garments from the inputs of clothed humans.
@@ 3.2 Representation of Base Pattern and Body
We assume that several base models, i.e., patterns and their corresponding sewn 3D meshes, are available for representative garment categories, which are set by the user depending on the target garment type under consideration. Our system provides three base patterns chosen from the Berkeley Garment Library [39], with the option for users to incorporate their own custom-created base models.
Figure 2: Control points on the T-shirt base pattern mesh. Specific seams are colored consistently across different panels for clarity.
Parameterization. The planar pattern mesh $U$ serves as the FEM reference (prior to any deformation) for its corresponding 3D garment mesh $X$ during the later simulation. The mapping between a 2D vertex $u\in U$ to a 3D node $x\in X$ is known from the base models, in the form of a UV Map. A panel is a triangular mesh bounded by a number of piece-wise curves, each composed of a number of edges. Two curves join at a control point $c^{i}$, which is typically the vertex of curvature discontinuity (corner point), or the vertex having three or more seam-counterparts belonging to other panels (join point) which will be merged into one node at the time of sewing, as shown in Fig. 2. The control points are grouped into disjoint sets $\mathcal{C}=\{C_{p}\}$, one for each panel. The points are in counterclockwise order within each panel $p$, i.e. $C_{p}=\{c_{p}^{i}\}$. Throughout the pattern optimization process, the control points serve as variables, while the remaining points will be repositioned so as to retain their relative locations with respect to them.
Symmetry detection. To further reduce the dimension of parameter space, and to preserve the left-to-right symmetry (which is often a desirable property in garment production) during optimization, we detect the pattern symmetry in two steps: It first detects inter-panel symmetry by computing for each pair of panels the aligning rigid transformation [47] and evaluating the quality of alignment. If their alignment score is sufficiently high, we remove one of them from the effective control point group ${K}$. Next, we perform intra-panel symmetry detection within ${K}$. It is done by computing for each pair of control points $c_{p}^{i},c_{p}^{j}\in C_{p}$ its axis of symmetry and evaluating the symmetry score for the remainder of control points in $p$. The control points pair with the highest score exceeding a predefined threshold will be used to identify left-to-right symmetry within a panel, which is then used to further reduce control points from ${K}$. The details of the symmetry detection method can be found in the supplementary material.
@@ 3.3 Pattern Linear Grading
In this phase, we aim to perform preliminary geometric deformation at the panel level to approximately capture the target garment by using a number of key measurements in 3D. The main idea is to focus on the open contours (i.e., 3D closed curves composed of edges connected to only one adjacent triangle) extracted from both the target and simulated meshes, and relocate pattern points so that the corresponding contours are closely aligned in 3D, in terms of their circumferences and their location along the bone. Note that open contours often carry design features, representing elements like neck lines, hem contours, cuff contours, etc.
Given a base model pair $U_{base}$ and $X_{base}$, an initial draped shape $X_{init}$ is computed on the estimated body $B$, with reference to $U_{base}$. The open contour curves on both the simulated mesh and the target mesh are extracted, associated with their respective counterparts, and the distances between them are measured along the skeleton of the underlying body. This distance, along with the difference in circumference, is used to guide the analytical relocation of control points $\{c^{i}\}$ on the pattern. An algorithmic description is provided in the supplementary material.
@@ 3.4 Cloth Simulation
The pattern obtained from the previous phase is only an approximation of the target geometry. In the next phase, we conduct pattern refinements through an optimization tightly coupled with a differentiable cloth simulation. In particular, we extend the differentiable ARCSim [32, 39] (Sect. 3.4.1-3.4.2) by revisiting both the dynamic solve (Sect. 3.4.3) and body-cloth interaction (Sect. 3.4.4).
@@ 3.4.2 Material Model
$\boldsymbol{H}$ $=\left[\begin{array}[]{lll}H_{00}&H_{01}&0\\ H_{01}&H_{11}&0\\ 0&0&H_{22}\end{array}\right].$ (2)
@@ 3.4.3 Acceleration of Force Vector Assembly
ARCSim [39, 32] uses the traditional approach of directly solving the linear system after the assembly of the extended mass matrix $\hat{\mathbf{M}}$ and the force vector $\hat{\mathbf{f}}$. The internal forces exerted by a triangle element to its nodes are split and accumulated to the global force vector, where the contributions from multiple adjacent elements are summed up for each node. Such force vector assembly process incurs a considerable overhead cost, as the number of time steps grows. It’s even more expensive for extended mass matrix assembly as it contains the Jacobian of forces, which is large and sparse. We propose an efficient method for accelerating the assembly. As the same assembly is executed for each time integration, we exploit the fact that the inherent topological structure remains unchanged during the simulation, with a sequence of triangle elements in a fixed filling order. We encode this information in a form of a static mapping matrix which maps the batched elementary forces/jacobians to the global force vector/jacobian matrix. This mapping converts the assembly process to a matrix multiplication as depicted in Fig. 3, which is parallelized on a GPU. Additionally, the matrix remains very sparse regardless of the mesh resolution, for which multiple numerical tools are available. It is highly backpropagation-friendly, which boosts the speed of the optimization phase (Table 1). More details on this scheme are provided in the supplementary material.
@@ 3.5 Optimization-based Pattern Alteration
$\quad+\lambda_{seam}\mathcal{L}_{seam}(U(K)),$
where $\lambda$’s are weights. It combines the reconstruction loss $\mathcal{L}_{rec}$ and the seam-consistency loss $\mathcal{L}_{seam}$ penalizing the inconsistent curve lengths along the seam. The reconstruction error is composed of the Chamfer distances $\mathcal{L}_{CF}$ measured both between the surfaces ($X$ and $T$) and among open contours ($X_{open}$ and $T_{open}$), and the loss $\mathcal{L}_{mat}$ penalizing unrealistic material parameters:
$\mathcal{L}_{rec}$ $=\mathcal{L}_{CF}(X,T)+\lambda_{open}\mathcal{L}_{CF}(X_{open},T_{open})$ (6)
$\quad+\lambda_{mat}\mathcal{L}_{mat}(\Gamma).$
@@
$\mathcal{L}_{CWCF}\left(X,T\right)$ $=\frac{1}{\left|X\right|}\sum_{\mathbf{x}\in X}\kappa(\tilde{\mathbf{t}})\min_{\mathbf{t}\in T}\|\mathbf{x}-\mathbf{t}\|^{2}$ (7)
$+\frac{1}{\left|T\right|}\sum_{\mathbf{t}\in T}\kappa(\mathbf{t})\min_{\mathbf{x}\in X}\|\mathbf{t}-\mathbf{x}\|^{2},$
@@
where $\mathcal{S}=\{\mathcal{S}_{i}\}\subset X$ denotes the set of 3D seam curves, and $E_{i},E_{i}^{corr}\subset U$ are the sets of edges $e^{j}=u^{j}-u^{j+1}$ along the 2D panel contours comprising the seam counterparts of $\mathcal{S}_{i}$.
@@ 4.1 Implementation details
Simulation. We set one time step $\Delta t$ to 0.05s, and the number of time steps for one forward simulation between 10 and 20. The garment resolution tested ranges from 1K to 3K vertices. By vectorizing as much as possible the force and jacobian computation, our extension to ARCSim differentiable cloth simulator [32] allows to run all its computations on a GPU. Tab. 1 summarizes the computation time about the linear solve of the baseline model [32] and ours measured on a NVIDIA GeForce 3090, for T-shirt garments with different resolutions (1K to 3K). Note that in our model the first iteration involves the construction of the deterministic mapping, which is reused for the subsequent iterations. On the contrary, the overhead of assembly is required for every iterations in the baseline model. We compared our SDF query network with classical KDTree-based SDF, and the evaluation times were 0.108s and 0.0054s, respectively, for 3K query points. The periodic activations used significantly reduce the average prediction error of SDF values (from 1.19mm to 0.2mm), similarly to the results of [50]. Overall, we observed 15 times speedup compared to [32].
@@ 4.2 Quantitative and Qualitative Comparisons
Figure 5: Comparison of 3D garment reconstruction of our method with others [41, 17, 18, 29]. Our method notably reproduces faithful garment shapes, even accounting for intricate geometry details like wrinkles on large sleeves. Best viewed on screen zoomed-in.
Figure 6: 2D pattern estimation from 3D garment meshes. From top to bottom: The ground truth, our results, and the results from NeuralTailor[27].
@@
Garments
SMPLicit
TailorNet Drapenet ISP Ours NeuralTailor Ours
T-shirt
1.4 /-
0.331 / 0.081 0.689 / 0.129 0.297 / 0.094 0.112 / 0.049 10.70 / 0.13 9.53 / 0.09
Dress
3.2 / -
1.305 / 0.161 0.619 / 0.135 0.189 / 0.131 0.110 / 0.075 11.3 / 0.37 10.96 / 0.10
Shorts
1.3 / -
1.036 / 0.050 0.131 / 0.048 0.202 / 0.095 0.126 / 0.043  7.61 / 0.05 7.37 / 0.04
Pants
2.9 / -
2.587 / 0.104 0.485 / 0.085 0.185 / 0.077 0.142 / 0.049 6.89 / 0.01 7.12 / 0.08
Skirt
6.5 / -
1.30 / 0.063 - / - 0.435 / 0.093 0.106 / 0.014 4.54 / 1.11 3.99 / 0.04
@@ 4.4 Evaluation on 3D Scan Data and Retargeting
We conducted a qualitative evaluation of our method using a 3D body scan. As shown in Fig. 8, it outputs reasonable, quality estimations of the 3D garment and the 2D pattern.
Figure 8: Results of our method evaluated using 3D scan data. (a) Input 3D scan; (b) Segmented target (left) and simulated garments (right); (c) Ground-truth pattern; and (d) Estimated pattern.
We also showcase an example of how our recovered garment can be retargeted to new conditions. As the recovered pattern is simulation compatible, it can be easily reused by a simulator to generate new draping shapes. As Fig. 9 shows, the dress model recovered from the 3D scan (Fig. 8) has been successfully used to produce compelling results on novel body poses and shapes.
Figure 9: The dress model obtained from the 3D scan (Fig. 8) has been retargated to new SMPL bodies.
@@ 5 Conclusion
Supplementary Material
6 Base Pattern
As mentioned in Sect. 3.2, our system provides base models selected and modified from the Berkeley Garment Library [39], while allowing the user to introduce personalized or customized base models into the system. Fig. 10 shows three base models (T-shirt/dress, pants, and skirts) used in our experiments.
Figure 10: The template models for three base topologies. (a) T-shirt/dress topology; (b) pants topology; (c) skirt topology.
7 Matrix Assembly Acceleration
Similar to the force vector $\hat{\mathbf{f}}$, the matrix $\hat{\mathbf{M}}$ discussed in Sec. 3.4.1 is constructed by aggregating the local contributions from individual elements into the corresponding locations.  For each triangle element (element hereafter), a collection of nine Jacobians is computed, each one being a partial derivative of a force with respect to a node’s position ($\frac{\partial f_{i}}{\partial x_{j}}$), $i$,$j$=1,2,3. The Jacobians for all elements are packed into a Jacobian stack as shown in Fig. 11 (d). For a compact representation, we use $ij$ to denote $\frac{\partial f_{i}}{\partial x_{j}}$ in the diagram. In ARCSim [32, 39], these Jacobians are assembled to $\hat{\mathbf{M}}$ by means of $F\times 3\times 3$ ($F$: number of triangles in the garment mesh) number of assign or add operations, following the list of elements. Given that this process incurs an even higher computational cost compared to the assembly of force vector, optimizing it becomes crucial.
In this work, we propose to realize the Jacobian assembly through matrix multiplication. $\hat{\mathbf{M}}$ is a sparse matrix, unlike the global force vector. Furthermore, it is impossible to directly represent the mapping from the Jacobian stack to it in matrix form, which would be ideal for leveraging the GPU-accelerated matrix multiplication. We address these issues by introducing an intermediate data structure called compressed Jacobian vector (Fig. 11(a)). It is a set of Jacobians for each force-node combination, which is obtained by first reshaping the Jacobian stack into a Jacobian vector (Fig. 11(c)), and by encoding the mapping from the per-element Jacobian to the compressed Jacobian vector as a static mapping matrix (Fig. 11(b)). Then a GPU-based sparse matrix multiplication can be performed, effectively substituting the iteration-based Jacobian assembly. This results in a considerable acceleration of the assembly process.
Figure 11: Assembly of compressed Jacobian vector is obtained by static matrix multiplication, encoding the mapping from the per-element Jacobian to the per force-node Jacobian.
Finally, the Jacobians in the compressed vector are transferred to the $\hat{\mathbf{M}}$ (Fig.12). Note that the number of operations reduces to $N+2\times E$ ($N$: number of nodes, $E$: number of edges), compared to the original $F\times 3\times 3$.

### 15 Mar 2024 v2

@@ Abstract
Keywords:
Differentiable simulation Garment modeling Pattern recovery
@@ 1 Introduction
We evaluate our approach on a variety of garment types and demonstrate that our method produces patterns of promising quality. We compare it against the state-of-the-art methods, and report that our method can replicate more detailed deformations and produce more accurate pattern estimations. In summary, we make the following contributions:
@@ 2.1 Garment shape recovery from 2D/3D data
Our work is similar in spirit to that of [60], who use search-based optimization to recover both 2D sewing patterns (parameterized by numerical, such as sleeve length, waist width) and the 3D garment such that it minimizes the 2D silhouette difference between the projected 3D garment and the input image. The draped 3D garment shape is obtained by using a physics-based simulator [40]. Unlike theirs, in our work the estimation of pattern parameters are coupled with the physically-based simulation process, allowing for direct and precise mapping of 3D garment error to both the pattern geometry and material parameters.
@@ 2.2 Sewing pattern estimation from 3D data
3D-to-2D surface flattening. Several works considered the garment mesh as developable surface and obtained 2D pattern panels by cutting the garment surface into 3D patches and flattening each of them onto a plane. The cutting lines are found either by projecting the predefined seam lines from the body mesh onto the garment mesh [2], along curvature directions expressed in the cross fields [54, 43] on the surface, or through variational surface cutting[49] to minimize the distortion induced by cutting and flattening [58]. While intuitive and versatile, such surface flattening based on a geometric strategy is prone to generating patterns that deviate from traditional panel semantics, or lack of left-to-right symmetry, making them unsuitable for garment production. Moreover, purely geometric methods [2, 43] do not account for the fabric’s elasticity in physical body-cloth interaction during the draping process, often resulting in a sewing pattern that cannot accurately replicate the originally designed garment.
@@ 2.3 Differentiable cloth simulator
Our method mainly solves an optimization of parameters coupled with the garment simulation process. We follow the thread of differentiable physics simulation [32, 22, 30, 28, 19] that can optimize parameters to fit the measurement data. To our knowledge, this is the first work that applies the differentiable simulator to recover sewing patterns of garments from the inputs of clothed humans.
@@ 3.2 Representation of Base Pattern and Body
We assume that several base models, i.e., patterns and their corresponding sewn 3D meshes, are available for representative garment categories, which are set by the user depending on the target garment type under consideration. Our system provides three base patterns chosen from the Berkeley Garment Library [40], with the option for users to incorporate their custom-created base models.
Parameterization. The planar pattern mesh $U$ serves as the FEM reference (prior to any deformation) for its corresponding 3D garment mesh $X$ during the later simulation. The mapping between a 2D vertex $u\in U$ to a 3D node $x\in X$ is known from the base models, in the form of a UV Map. A panel is a 2D triangular
mesh bounded by a number of piece-wise curves, each composed of a number of edges. Two curves join at a control point $c^{i}$, which is typically the vertex of curvature discontinuity (corner point), or the vertex having three or more seam-counterparts belonging to other panels (join point) which will be merged into one node at the time of sewing.
The control points are grouped into disjoint sets $\mathcal{C}=\{C_{p}\}$, one for each panel. The points are ordered in a counterclockwise manner within each panel $p$, i.e. $C_{p}=\{c_{p}^{i}\}$. Throughout the pattern optimization process, the control points serve as variables, while the remaining points will be repositioned so as to retain their relative locations with respect to them.
Symmetry detection. To further reduce the dimension of parameter space, and to preserve the left-to-right symmetry (which is often a desirable property in garment production) during optimization, we detect the pattern symmetry in two steps: It first detects inter-panel symmetry by computing for each pair of panels the aligning rigid transformation [48] and evaluating the quality of alignment. If their alignment score is sufficiently high, we remove one of them from the effective control point group ${K}$. Next, we perform intra-panel symmetry detection within ${K}$, by computing for each pair of control points $c_{p}^{i},c_{p}^{j}\in C_{p}$ its axis of symmetry and evaluating the symmetry score for the remainder of control points in $p$. The control points pair with the highest score exceeding a predefined threshold will be used to identify left-to-right symmetry within a panel, which is then used to further reduce control points from ${K}$. The details of the symmetry detection method can be found in the supplementary material.
@@ 3.3 Pattern Linear Grading
In this phase, we aim to perform preliminary geometric deformation at the panel level to approximately capture the target garment by using a number of key measurements in 3D. The main idea is to focus on the open contours (i.e., 3D closed curves composed of edges connected to only one adjacent triangle) extracted from both the target and simulated meshes, and relocate control points so that the corresponding contours are closely aligned in 3D, in terms of circumferences and location along the bone. Note that open contours often carry design features, representing elements like neck lines, hem contours, cuff contours, etc.
Given a base model pair $U_{base}$ and $X_{base}$, an initial draped shape $X_{init}$ is computed on the estimated body $B$, with reference to $U_{base}$. The open contours on both the simulated and the target meshes are extracted, associated with their respective counterparts, the distances between them are measured along the skeleton of the underlying body. The distances, along with the difference in circumference, are used to guide the analytical relocation of control points $\{c^{i}\}$ on the pattern. An algorithmic description is given in the supplementary material.
@@ 3.4 Cloth Simulation
The pattern obtained from the previous phase is only an approximation of the target geometry. In the next phase, we conduct pattern refinements through an optimization tightly coupled with a differentiable cloth simulation. In particular, we extend the differentiable ARCSim [32, 40] by revisiting both the dynamic solve and body-cloth interaction.
@@ 3.4.2 Material Model
$\boldsymbol{H}$ $=\left[\begin{array}[]{lll}H_{00}&H_{01}&0\\ H_{01}&H_{11}&0\\ 0&0&H_{22}\end{array}\right],$ (2)
@@ 3.4.3 Acceleration of Force Vector Assembly
ARCSim [40, 32] uses the traditional approach of directly solving the linear system after the assembly of the extended mass matrix $\hat{\mathbf{M}}$ and the force vector $\hat{\mathbf{f}}$. The internal forces exerted by a triangle element to its nodes are split and accumulated to the global force vector, where the contributions from multiple adjacent elements are summed up for each node. Such force vector assembly process incurs a considerable overhead cost, as the number of time steps grows. It’s even more expensive for extended mass matrix assembly as it contains the Jacobian of forces, which is large and sparse. We propose an efficient method for accelerating the assembly. As the same assembly is executed for each time integration, we exploit the fact that the inherent topological structure remains unchanged during the simulation, with a sequence of triangle elements in a fixed filling order. We encode this information in a form of a static mapping matrix which maps the batched elementary forces/jacobians to the global force vector/jacobian matrix. This mapping converts the assembly process to a matrix multiplication as depicted in Fig. 2, which is parallelized on a GPU. Additionally, the matrix remains very sparse regardless of the mesh resolution, for which multiple numerical tools are available. It is highly backpropagation-friendly, which boosts the speed (Table 1). More details on this scheme are in the supplementary material.
@@ 3.5 Optimization-based Pattern Alteration
where $\lambda$’s are weights. It combines the reconstruction loss $\mathcal{L}_{rec}$ and the seam-consistency loss $\mathcal{L}_{seam}$ penalizing the inconsistent curve lengths along the seam. The reconstruction error is composed of the Chamfer distances $\mathcal{L}_{CF}$ measured both between the surfaces ($X$ and $T$) and among open contours ($X_{open}$ and $T_{open}$),
$\mathcal{L}_{rec}=\mathcal{L}_{CF}(X,T)+\lambda_{open}\mathcal{L}_{CF}(X_{open},T_{open})+\lambda_{mat}\mathcal{L}_{mat}(\Gamma).$ (6)
@@
$\mathcal{L}_{CWCF}\left(X,T\right)=\frac{1}{\left|X\right|}\sum_{\mathbf{x}\in X}\kappa(\tilde{\mathbf{t}})\min_{\mathbf{t}\in T}\|\mathbf{x}-\mathbf{t}\|^{2}+\frac{1}{\left|T\right|}\sum_{\mathbf{t}\in T}\kappa(\mathbf{t})\min_{\mathbf{x}\in X}\|\mathbf{t}-\mathbf{x}\|^{2},$ (7)
@@
where $\mathcal{S}=\{\mathcal{S}_{i}\}\subset X$ denotes the set of 3D seam curves, and $E_{i},E_{i}^{corr}\subset U$ are the sets of edges $e^{j}=u^{j}-u^{j+1}$ along the 2D panel curves comprising the seam counterparts of $\mathcal{S}_{i}$.
@@
Figure 4: Up: Comparison of 3D garment reconstruction of our method with others [42, 16, 17, 29]. Our method reproduces faithful garment shapes, even accounting for intricate geometry details like wrinkles on large sleeves. Down: 2D pattern estimation from 3D garment meshes. From top to bottom: The ground truth, our results, and the results from NeuralTailor[27]. Best viewed on screen zoomed-in.
@@ 4.1 Implementation details
Simulation. We set one time step $\Delta t$ to 0.05s, and the number of time steps for one forward simulation between 10 and 20. The garment resolution tested ranges from 1K to 3K vertices. By vectorizing as much as possible the force and jacobian computation, our extension to ARCSim differentiable cloth simulator [32] allows to run all its computations on a GPU. Table 1 summarizes the computation time of linear solve for the baseline model [32] and ours measured on a NVIDIA GeForce 3090, for T-shirt garments with different resolutions (1K to 3K). Note that in our model the first iteration involves the construction of the deterministic mapping, which is reused for the subsequent iterations. On the contrary, the overhead of assembly is required for every iterations in the baseline model. We compared our SDF query network with classical KDTree-based SDF, and the evaluation times were 0.108s and 0.0054s, respectively, for 3K query points. The periodic activations used significantly reduce the average prediction error of SDF values (from 1.19mm to 0.2mm), similarly to the results of [51]. Overall, we observed 15 times speedup compared to [32].
@@ 4.2 Quantitative and Qualitative Comparisons
Garments SMPLicit TailorNet Drapenet ISP Ours NeuralTailor Ours
T-shirt 1.4 /- 0.331 / 0.081 0.689 / 0.129 0.297 / 0.094 0.112 / 0.049 10.70 / 0.13 9.53 / 0.09
Dress 3.2 / - 1.305 / 0.161 0.619 / 0.135 0.189 / 0.131 0.110 / 0.075 11.3 / 0.37 10.96 / 0.10
Shorts 1.3 / - 1.036 / 0.050 0.131 / 0.048 0.202 / 0.095 0.126 / 0.043  7.61 / 0.05 7.37 / 0.04
Pants 2.9 / - 2.587 / 0.104 0.485 / 0.085 0.185 / 0.077 0.142 / 0.049 6.89 / 0.01 7.12 / 0.08
Skirt 6.5 / - 1.30 / 0.063 - / - 0.435 / 0.093 0.106 / 0.014 4.54 / 1.11 3.99 / 0.04
@@ 4.3 Recovery of Physical Parameters
Figure 6: Results of our method evaluated using 3D scan data. (a) Input 3D scan; (b) Segmented target (left) and simulated garment (right); (c) Ground-truth pattern; (d) Estimated pattern; (e) Retargeted garment to two new SMPL bodies.
@@ 4.4 Evaluation on 3D Scan Data and Retargeting
We conducted a qualitative evaluation of our method using a 3D body scan. As shown in Fig. 6, it outputs reasonable, quality estimations of the 3D garment and the 2D pattern. We also showcase an example of how our recovered garment can be retargeted to new conditions. As the recovered pattern is simulation compatible, it can be easily reused by a simulator to generate new draping shapes. The dress model recovered from the 3D scan has been successfully used to produce compelling results on novel body poses and shapes.
@@ 5 Conclusion
Appendix 0.A Base patterns
Our system provides base models selected and modified from the Berkeley Garment Library [40] (Sect. 3.2 in our paper), which users can modify as needed. Additionally, users have the option to import their own personalized or customized base models into the system. Fig. 7 shows three base models (t-shirt/dress, pants, and skirt) used in our experiments.
Figure 7: The base models for three garment topologies. (a) t-shirt/dress; (b) pants; and(c) skirt.
Appendix 0.B Matrix assembly acceleration
Similar to the force vector $\hat{\mathbf{f}}$, the matrix $\hat{\mathbf{M}}$ in Sect. 3.4 of our paper is constructed by aggregating the local contributions from individual elements into the corresponding locations.  Each triangle element (element hereafter) yields nine Jacobians, each representing a partial derivative of the force with respect to the position of a node ($\frac{\partial f_{i}}{\partial x_{j}}$), $i,j$=1,2,3. The Jacobians for all elements are packed into a Jacobian stack as illustrated in Fig. 8 (d), where we use $mn$ to denote $\frac{\partial f_{m}}{\partial x_{n}}$ ($m$, $n$: global indices) for a compact representation. In ARCSim [32, 40], these Jacobians are assembled to $\hat{\mathbf{M}}$ through a total of $F\times 3\times 3$ assignment or addition operations, where $F$ represents the number of triangles in the garment mesh. Optimizing this process becomes crucial, especially considering its higher computational cost compared to the force vector assembly (The optimization of the latter has been discussed in Sect. 3.4 of the main paper).
To this end, we propose to realize the Jacobian assembly using matrix multiplication. $\hat{\mathbf{M}}$ is a sparse matrix, and moreover, directly representing the mapping from the Jacobian stack to the matrix form is not feasible, although it would be ideal for leveraging GPU-accelerated matrix multiplication. We address this issue by introducing an intermediate data structure called compressed Jacobian vector (Fig. 8(a)). It is a set of Jacobians for each force-node combination, which is obtained by first reshaping the Jacobian stack into a Jacobian vector (Fig. 8(c)), and by encoding the mapping from the per-element Jacobian to the compressed Jacobian vector as a static mapping matrix (Fig. 8(b)). Then a GPU-based sparse matrix multiplication is performed, effectively substituting the iteration-based Jacobian assembly. This results in a considerable acceleration of the assembly process as reported in Table. 1 of our paper.
Figure 8: The assembly of a compressed Jacobian vector is obtained by a static matrix multiplication, encoding the mapping from the per-element Jacobian (d) to the per force-node Jacobian (a).
Finally, the Jacobians in the compressed vector are transferred to $\hat{\mathbf{M}}$ (Fig.9). Note that the number of operations reduces to $N+2\times E$ ($N$: number of nodes, $E$: number of edges), compared to the original $F\times 3\times 3$.